OpenStudy (anonymous):

how do you write 3.21 with the dash over the 21 as a mixed number

OpenStudy (anonymous):

so 3.21(repeating)?

OpenStudy (anonymous):

just the 21

OpenStudy (anonymous):

mabe @just_one_last_goodbye or @magan @greenglasses @Alchemista @Vcarias123 @iamabarbiegirl @nincompoop @One098 @willreel @countrygirl1431

OpenStudy (anonymous):

3 and 190/9

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

if theres a dash over 21 its a repeating decimal

OpenStudy (anonymous):

right

OpenStudy (unklerhaukus):

x = 3.21212121... 100x = 321.21212121... 99x =

OpenStudy (anonymous):

so 3 and 190/9 is what I have

OpenStudy (anonymous):

\[x=3.21212121\]\[100x=321.212121\]\[100x-x=99x=321.212121-3.212121=318\]\[9x=318\]\[x=318/9\]

OpenStudy (anonymous):

Made a typo in the last two lines: \[99x = 318\]\[x=\frac{318}{99}\]

OpenStudy (unklerhaukus):

careful

OpenStudy (anonymous):

oops I realized 21 was repeating not just the .01

OpenStudy (anonymous):

thats not the correct form it has to be something like 2 3/4

OpenStudy (anonymous):

Can you convert 318/99 into a mixed fraction? I've done 99% of the work for you.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

We know ac+b=318 and that c=99, so you can work out b: \[a\frac{b}{c} = \frac{ac+b}{c}\]

OpenStudy (anonymous):

i still don't get it

OpenStudy (anonymous):

This is very basic maths, if you are being asked to find fractions from recurring decimals you should be able to convert an improper fraction to a mixed fraction. I'll do another fraction as an example: \[\frac{32}{7} = \frac{4*7+3}{7} = \frac{ 4*7}{7}+\frac{3}{7} = 4 \frac{3}{7}\] Find how many times 99 (your c) goes into 318 (this is your a), then find the remainder (your b).